OpenStudy (anonymous):

in a simple model of the hydrogen atom, the electron moves in a circular orbit of radiu 0.053nm around a stationary proton. how many revs/sec does the electron make?

OpenStudy (anonymous):

don't i need speed?

OpenStudy (jamesj):

yes you do

OpenStudy (jamesj):

...unless you've discussed in class other parameters of this model of an atom which enable you to calculate speed from what you've been given. But for what it's worth, I don't know it and it's just not standard, because this is not the modern model of the atom.

OpenStudy (anonymous):

If you are thinking of Bohr model of the atom then you can think in terms of forces acting on the electron to get the velocity. You can say that electron is held in a circular orbit by Coulomb force and that this force is equal to the centripetal force:\[\ F_{centripetal} = F_{Coulomb}\] \[\ \frac{m_ev^2}{r} = \frac{Zk_ee^2}{r^2} \] where: m_e - mass of electron v- velocity of the electron r - radius of the orbit z - atomic number \[\ k_e=\frac{1}{4\pi \epsilon_0} \] e - charge of electron Solve this equation for v to get the velocity of electron in orbit with radius r Also note that Bohr model is incorrect and it's not a modern view of the atom, like JamesJ said. I guess they still teach it so the students can get some "feel" for the subject.

OpenStudy (jamesj):

yes, gogind is exactly right. I should have thought of that.

OpenStudy (anonymous):

what does \[\epsilon _{0}\] represent?

OpenStudy (jamesj):

it's a constant known as the permittivity of free space. It's a constant in the Coulomb force equation.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

what a messsss. if you have the time, will you tell me what you got? i got 2.18*10^6...i guess m/s. not sure how to convert the units to m/s but that's the # i got.

OpenStudy (anonymous):

oops, that's the velocity. still haven't got revs/sec...

OpenStudy (anonymous):

1.53*10^-4....doesn't seem right since it would seem like it should be a big #

OpenStudy (anonymous):

nope...FINAL ANSWER=4.11*10^16 rev/sec